Dinitrogen oxide, commonly called laughing gas, is sometimes used as an anesthetic in dentistry. It is produced by the decomposition of ammonium nitrate. The other product of this reaction is water. A balanced equation indicates the mole ratios of reactants and products. The molar mass of a substance allows you to convert between moles and mass.

______________________________________________________________________________

Write the balanced equation for the production of dinitrogen oxide and water from ammonium nitrate.

How many grams of ammonium nitrate are required to produce 66.0g of dinitrogen oxide.

How many grams of water are produced by this reaction, in consideration of 66.0g N2O?

How many grams of laughing gas would be produced from the decomposition of 60g of ammonium nitrate? (read paragraph above)
Answer #1

The reaction that occurs in:

NH4NO3 = N2O + 2 H2O

The mass of NH4NO3 required is calculated:

m NH4NO3 = 66 g N2O * (1 mol N2O / 44 g) * (1 mol NH4NO3 / 1 mol N2O) * (80 g NH4NO3 / 1 mol) = 120 g

The mass of water produced is calculated:

m H2O = 66 g N2O * (1 mol N2O / 44 g) * (2 mol H2O / 1 mol N2O) * (18 g H2O / 1 mol) = 54 g

The mass of N2O produced is calculated:

m N2O = 60 g NH4NO3 * (1 mol NH4NO3 / 80 g) * (1 mol N2O / 1 mol NH4NO3) * (44 g N2O / 1 mol) = 33 g

If you liked the answer, please rate it positively, you would help me a lot, thanks.